OpenStudy (anonymous):

i need help in algebra anyone willing to help?(:

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Simplify \[\left( 2\sqrt{5+3\sqrt{7^2}} \right)\] can someone show me step by step?

OpenStudy (anonymous):

What is : \[\sqrt{7^2}= ??\]

OpenStudy (anonymous):

Hey @Hoa you continue, I think I have disturbed you mid where..

OpenStudy (anonymous):

yes i know can you just try to help solve it solve what you think its suppose to be please

OpenStudy (anonymous):

I can't solve it on my own, you have to co operate otherwise it is not worth learning here...

OpenStudy (anonymous):

well ill co operatre

OpenStudy (anonymous):

Then tell me : \[\sqrt{7^2} = ??\]

OpenStudy (anonymous):

Take it apart piece by piece, first dealing with the innermost term \[\sqrt{7^{2}} =7\] Then look at the next part:\[\sqrt{5+(3*7)}=\sqrt(5+21)=\sqrt(26)\] And then the final part: \[2\sqrt(26)\] Just go with the idea of taking the innermost terms and slowly expanding it until you have looked at the entire equation.

OpenStudy (anonymous):

thanks :)